# A Survey on Large Language Models with Multilingualism: Recent Advances and New Frontiers

### v1

@@ 2.3 Multilingual Paradigm Transition
As shown in Figure 2, multilingual pre-trained language models (PLMs) are dominated by small language models (i.e., parameters less than 7B). They are designed for specific tasks, such as multilingual neural machine translation (NMT) Johnson et al. (2017); Aharoni et al. (2019); Firat et al. (2017), multilingual question answering (QA) Loginova et al. (2018, 2021); Ruder and Sil (2021) and multilingual reasoning Lin et al. (2021b); Fang et al. (2022); Bang et al. (2023), etc. To indicate the language of target tasks, a prepending language token is appended to each source sentence Johnson et al. (2017); Conneau and Lample (2019). Take the multilingual NMT for example, formally, given the source sentence $\mathbf{x^{\prime}}=(x_{1},x_{2},...,x_{n})$, where each $x_{i}$ is a token in the source language, the modified source sentence is represented as $\mathbf{x}=(l_{i},x_{1},x_{2},...,x_{n})$. Here, $l_{i}$ denotes the target language. Correspondingly, the target sentence is represented as $\mathbf{y}=(y_{1},y_{2},...,y_{m})$, where each $y_{j}$ represents a token in the target language. The probability of a target sentence is given by:
@@
However, the current model paradigm has gradually shifted to Artificial General Intelligence (AGI) which possesses the ability like human intelligence to perform a wide range of tasks Ouyang et al. (2022); Achiam et al. (2023); Yang et al. (2023b), including understanding, learning, cognition, communication and others. Unlike small language models, which are designed for specific tasks, AGI would have the capacity for generalization and adaptation to novel situations, potentially surpassing human capabilities in various domains. With the increasing scale of language models, LLMs improve the performance of downstream tasks compared to small language models Kaplan et al. (2020), which is closer to the goal of AGI. For the sake of generality, the task form becomes general tasks instead of specific tasks, and prepending the language is not allowed. Hence, most existing LLMs typically do not specify whether they function as multilingual models or detail the number of languages they support. We consider that if a certain amount of multilingual data is used in the training corpus, it can be regarded as a multilingual LLM. Among the existing models, only a few are explicitly called “multilingual LLM” (e.g., BLOOM Le Scao et al. (2022), BayLing Zhang et al. (2023a), etc), yet there are also other models that possess multilingual capabilities Jiang et al. (2023); Wang et al. (2023c). To distinguish LLMs from multilingual LLMs, we propose a definition of “LLMs with multilingualism”.
{forest}
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center, calign=edge midpoint, , [LLMs with Multilingualism, text width=1.7cm, fill=black!10 [Representative Models §3, text width=1.8cm, for tree=fill=red!20 [Training from scratch §3.1, text width=1.8cm [ GPT-3 Brown et al. (2020); mT5 Xue et al. (2020); ByT5 Xue et al. (2022); Gopher Rae et al. (2021); LaMDA Thoppilan et al. (2022); OPT Zhang et al. (2022a); PaLM Chowdhery et al. (2023); mGPT Shliazhko et al. (2022); GPT-3.5 Ouyang et al. (2022); XGLM Lin et al. (2022); LLaMA Touvron et al. (2023); GPT-4 Achiam et al. (2023); PANGU-$\sum$ Ren et al. (2023); Pythia Biderman et al. (2023); PaLM-2 Anil et al. (2023); InternLM Team (2023); PolyLM Wei et al. (2023a); LLaMA-2 Touvron et al. (2023); Baichuan-2 Yang et al. (2023c); Qwen Bai et al. (2023); Mistral Jiang et al. (2023); Gemini Team et al. (2023); TigerBot Chen et al. (2023a); YAYI-2 Luo et al. (2023b); DeepSeek Bi et al. (2024); Orion Chen et al. (2024a); TeleChat Abdin et al. (2024); Claude3 Anthropic (2024); InternLM2 Cai et al. (2024); Phi-3 Abdin et al. (2024); LLaMA-3 AI@Meta (2024), text width=7.0cm, node options=align=left ] ] [Continual Training §3.2, text width=1.8cm [ FLAN-T5 Chung et al. (2024); FLAN-PaLM Chung et al. (2024); ChatGLM Zeng et al. (2022); Alpaca Taori et al. (2023); ParroT Jiao et al. (2023); BigTrans Yang et al. (2023a); Vicuna Chiang et al. (2023); BayLing Zhang et al. (2023a); OpenChat Wang et al. (2023c); GLM-4 ZHIPU (2024), text width=7.0cm, node options=align=left ] ] ] [Inference Strategies §4, text width=1.8cm, for tree=fill=yellow!20 [Translation-Based, text width=1.8cm [ Intrator et al. Intrator et al. (2024); Liu et al. Liu et al. (2024a), text width=7.0cm, node options=align=left ] ] [Chain-of-Thought §4.3, text width=1.8cm [ Liu et al. Liu et al. (2024a); XLT Huang et al. (2023c); Shi et al. Shi et al. (2022a); Kim et al. Kim et al. (2023a); Suzgun et al. Suzgun et al. (2023); xCoT Chai et al. (2024), text width=7.0cm, node options=align=left ] ] [Retrieval Augmented Generation §4.4, text width=1.8cm [ Zhang et al. Zhang et al. (2023b); Shi et al. Shi et al. (2022b); Agrawal et al. Agrawal et al. (2023); Li et al. Li et al. (2023a); Li et al. Li et al. (2023b); Winata et al. Winata et al. (2023a); Garcia et al. Garcia et al. (2023); Ramos et al. Ramos et al. (2023); Kim et al. Kim et al. (2023b); Thakur et al. Thakur et al. (2024); Sennrich et al. Sennrich et al. (2023); Vernikos et al. Vernikos and Popescu-Belis (2024); Fu et al. Fu et al. (2024); Zeng et al. Zeng et al. (2024); He et al. He et al. (2024); Conia et al. Conia et al. (2023), text width=7.0cm, node options=align=left ] ] [Code-Switching §4.5, text width=1.8cm [ Zhang et al. Zhang et al. (2023c); Koto et al. Koto et al. (2024); Peng et al. Peng et al. (2023a), text width=7.0cm, node options=align=left ] ] ] [Security §5, text width=1.8cm, for tree=fill=blue!20 [Attack (Jailbreak) §5.1, text width=1.0cm, for tree=fill=blue!20 [Greedy Coordinate Gradient, text width=1.4cm [ Sitawarin et al. Sitawarin et al. (2024); Zou et al. Zou et al. (2023), text width=6.0cm, node options=align=left ] ] [Prompt-Based, text width=1.4cm [ Wei et al. Wei et al. (2024); Liu et al. Liu et al. (2023b); Shen et al. Shen et al. (2023); Deng et al. Deng et al. (2024); Li et al. Li et al. (2024b); Liu et al. Liu et al. (2023c); Jin et al. Jin et al. (2024), text width=6.0cm, node options=align=left ] ] [Multilingual, text width=1.4cm [ Shen et al. Shen et al. (2024); Deng et al. Deng et al. (2023); Puttaparthi et al. Puttaparthi et al. (2023); Yong et al. Yong et al. (2023); Xu et al. Xu et al. (2023); Li et al. Li et al. (2024c); Yuan et al. Yuan et al. (2024); Huang et al. Huang et al. (2023d), text width=6.0cm, node options=align=left ] ] ] [Defense §5.2, text width=1.0cm, for tree=fill=blue!20 [Open-Source Models, text width=1.4cm [ Robey et al. Robey et al. (2023); Deng et al. Deng et al. (2023); Li et al. Li et al. (2024c); Zhou et al. Zhou et al. (2024a); GUARD Jin et al. (2024), text width=6.0cm, node options=align=left ] ] [Closed-Source Models, text width=1.4cm [ Jain et al. Jain et al. (2023); Wu et al. Wu et al. (2024); Li et al. Li et al. (2023c), text width=6.0cm, node options=align=left ] ] ] ] [Multidomain §6, text width=1.8cm, for tree=fill=green!20 [Medical §6.1, text width=1.8cm [ KBioXLM Geng et al. (2023); BioMistral Labrak et al. (2024); MMedLM2 Qiu et al. (2024a); Apollo Wang et al. (2024a); L2M3 Gangavarapu (2024); Medical mT5 Garc´ıa-Ferrero et al. (2024), text width=7.0cm, node options=align=left ] ] [Legal §6.2, text width=1.8cm [ LEXTREME Niklaus et al. (2023a); Brugger et al. Brugger et al. (2023); Christen et al. Christen et al. (2023); Baumgartner et al. Baumgartner et al. (2024); Niklaus et al. Niklaus et al. (2023b); LegalLAMA Chalkidis et al. (2023); Trautmann et al. Trautmann et al. (2022), text width=7.0cm, node options=align=left ] ] ] [Datasets, text width=1.8cm, for tree=fill=pink!20 [Corpus §7, text width=1.8cm [ Amazon intent mteb (a); Amazon reviews Keung et al. (2020a) ; Bactrian-x Li et al. (2023d); Biblenlp eBible.org ; Bloom-lm Le Scao et al. (2022); CC100 Conneau et al. (2020a) Wenzek et al. (2020); CulturaX Nguyen et al. (2023); GPT-4 Prompts erfanzar ; Guanaco Dettmers et al. (2023); HPLT de Gibert et al. (2024) ; IWSLT 2017 Cettolo et al. (2017); mC4 Raffel et al. (2019); Mewsli-x Ruder et al. (2021a); Minds14 Gerz et al. (2021) ; Miracl Zhang et al. (2022b) ; MLDR Chen et al. (2024b); MMedC Qiu et al. (2024b); MQA De Bruyn et al. (2021); Multi-sentiments tyqiangz ; Multiconer Fetahu et al. (2023a) Fetahu et al. (2023b); Nomiracl Thakur et al. (2023) Thakur et al. (2024);Open Subtitles Lison and Tiedemann (2016) ; OSCAR Abadji et al. (2022); Para-pat Soares et al. (2020) ; Project Gutenberg Gutenberg ; ShareGPT RyokoAI ; SREDFM Huguet Cabot et al. (2023); TED Talks Qi et al. (2018); TED-talks-iwslt Cettolo et al. (2012); Toxi-text FredZhang ; UD Nivre et al. (2020); Wikiann Pan et al. (2017) Rahimi et al. (2019); Wikipedia Foundation ; Wit Base Srinivasan et al. (2021); xP3 Muennighoff et al. (2022a), text width=7.0cm, node options=align=left ] ] [Benchmark §8, text width=1.8cm [ Afrisent Muhammad et al. (2023); ASPEN Razumovskaia et al. (2024) ; BELEBELE Bandarkar et al. (2023); BioInstructQA BioMistral ; Bucc-bitext-mining mteb (b); Cross-Sum Bhattacharjee et al. (2023); Crossmodal-3600 Thapliyal et al. (2022); Exams Hardalov et al. (2020); Fairlex Chalkidis et al. (2022a); FLORES-200 Team (2022) Goyal et al. (2022) Guzmán et al. (2019); GEOMLAMA Yin et al. (2022); Humaneval-XL Peng et al. (2024); M-Hellaswag Dac Lai et al. (2023); M-MMLU Dac Lai et al. (2023); M3Exam Zhang et al. (2023d); M3LS Verma et al. (2023); MARC Keung et al. (2020b); MasakhaNER Adelani et al. (2021); Masakhanews Adelani et al. (2023a); MASSIVE FitzGerald et al. (2022); MaXM Changpinyo et al. (2023); MEGA Ahuja et al. (2023); MEGAVerse Ahuja et al. (2024); Mela Zhang et al. (2024b); MGSM Shi et al. (2022a); MLQA Lewis et al. (2020b); MMedBench Qiu et al. (2024c); Multi-CoNER Malmasi et al. (2022); MULTIEURLEXDOC Chalkidis et al. (2021); Multilingual-Fig-QA Kabra et al. (2023); NusaX Winata et al. (2023b); ODEX Wang et al. (2023d); OPUS-100 Zhang et al. (2020); Paws-X Yang et al. (2019); Pmindiasum Urlana et al. (2023); PRESTO Goel et al. (2023); SEAHORSE Clark et al. (2023); Sib200 Adelani et al. (2023b); SMiLER Seganti et al. (2021); STSB-multi-mt May (2021); Tatoeba-mt Tiedemann (2020); Tydip Srinivasan and Choi (2022); TyDiQA Clark et al. (2020); Universal Dependencies Nivre et al. (2020) ; X-CLAIM Mittal et al. (2023); X-RiSAWOZ Moradshahi et al. (2023); XCOPA Ponti et al. (2020); XCSQA Lin et al. (2021b); xDial-Eval Zhang et al. (2023e); XGLUE Liang et al. (2020); XL-SUM Hasan et al. (2021); XNLI Conneau et al. (2018); XQuAD Artetxe et al. (2020); XSEMPLR Zhang et al. (2023f); XStoryCloze Lin et al. (2022); XTREME Hu et al. (2020); XTREME-R Ruder et al. (2021b); XWinograd Tikhonov and Ryabinin (2021); XCSR Lin et al. (2021c), text width=7.0cm, node options=align=left ] ] ] ]
@@ 3 Large Language Models with Multilingual Capability
GPT-3 Brown et al. (2020) 20-05 13B, 175B OpenAI GPT-2 Closed -
mT5 Xue et al. (2020) 20-10 13B Google T5 Open 101
ByT5 Xue et al. (2022) 21-05 13B Google T5 Open 101
Gopher Rae et al. (2021) 21-12 280B DeepMind - Open -
LaMDA Thoppilan et al. (2022) 22-01 137B Google - Open -
OPT Zhang et al. (2022a) 22-04 175B Meta - Open -
PaLM Chowdhery et al. (2023) 22-04 8B, 62B, 540B Google - Closed 124
mGPT Shliazhko et al. (2022) 22-04 13B - GPT-3 Open 61
BLOOM Le Scao et al. (2022) 22-07 176B BigScience - Open 46
FLAN-T5 Chung et al. (2024) 22-10 11B Google T5 Open 60
FLAN-PaLM Chung et al. (2024) 22-10 8B, 62B, 540B Google PaLM Open 60
ChatGLM Zeng et al. (2022) 22-10 130B ZHIPU GLM Open 2
GPT-3.5 Ouyang et al. (2022) 22-11 - OpenAI GPT Closed -
XGLM Lin et al. (2022) 22-11 7.5B Meta AI - Open 30
LLaMA Touvron et al. (2023) 23-02 7B, 13B, 33B, 65B Meta - Open -
GPT-4 Achiam et al. (2023) 23-03 - OpenAI - Closed -
Alpaca Taori et al. (2023) 23-03 7B StandFord LLaMA Open -
PANGU-$\sum$ Ren et al. (2023) 23-03 1085B Huawei - Closed 26
Pythia Biderman et al. (2023) 23-04 12B EleutherAI - Open -
ParroT Jiao et al. (2023) 23-04 7B Tencent Alpaca Open -
BigTrans Yang et al. (2023a) 23-05 13B - LLaMA Open 102
PaLM-2 Anil et al. (2023) 23-05 340B Google PaLM Closed -
Vicuna Chiang et al. (2023) 23-06 13B LMSYS LLaMA Open -
InternLM Team (2023) 23-06 104B Shanghai AI Laboratory - Open -
BayLing Zhang et al. (2023a) 23-06 7B, 13B ICT/CAS - Open -
PolyLM Wei et al. (2023a) 23-07 13B DAMO Academy - Open 18
LLaMA-2 Touvron et al. (2023) 23-07 7B, 13B, 34B, 70B Meta LLaMA Open -
Baichuan-2 Yang et al. (2023c) 23-09 7B, 13B Baichuan - Open -
Qwen Bai et al. (2023) 23-09 7B, 14B Alibaba - Open -
OpenChat Wang et al. (2023c) 23-09 7B, 13B Tsinghua LLaMA Open -
Mistral Jiang et al. (2023) 23-10 7B Mistral - Open -
Gemini Team et al. (2023) 23-12 - DeepMind - Closed 40+
TigerBot Chen et al. (2023a) 23-12 7B, 13B, 70B, 180B Tiger - Open -
YAYI-2 Luo et al. (2023b) 23-12 30B IACAS - Open -
DeepSeek Bi et al. (2024) 24-01 67B DeepSeek AI - Open -
GLM-4 ZHIPU (2024) 24-01 - ZHIPU GLM Close -
Orion Chen et al. (2024a) 24-01 14B OrionStar - Open 8+
TeleChat Abdin et al. (2024) 24-01 7B, 12B China Telecom - Open -
Claude3 Anthropic (2024) 24-03 - Anthropic - Closed -
InternLM-2 Cai et al. (2024) 24-03 7B, 20B Shanghai AI Lab Open -
LLaMA-3 AI@Meta (2024) 24-04 8B, 70B Meta - Open 30+
Phi-3 Abdin et al. (2024) 24-04 7B, 14B Microsoft - Open -
@@ 3.1 Training from Scratch
To obtain a language model with multilingual capability, a common practice is to leverage all available data in different languages for training. The language sampling algorithms are usually applied to control the importance of each language. As shown in Figure 2, multilingual pre-trained language PLMs before 2022 mainly consisted of two structures: (1) encoder-only designed for natural language understanding tasks and (2) encoder-decoder designed for natural language generation tasks. The parameters of these PLMs are not comparable with the existing LLMs. Recent studies show that according to the scaling law Kaplan et al. (2020); Muennighoff et al. (2024); Aghajanyan et al. (2023), the scale of the parameter has significant impacts on the performance of the models, i.e., the larger models lead to better performance. The researchers also observe that when the parameter scale of the language models exceeds a certain level, they not only achieve significant performance improvements but also find some special capabilities not observed within small-scale PLMs Kaplan et al. (2020); Muennighoff et al. (2024). To distinguish the difference in parameter scale, the research community specifies the term “LLMs” compared with PLMs based on a significant scale Zhao et al. (2023).
@@
In addition to constructing multilingual data in the pre-training stage, the cross-lingual capabilities can also be improved by some specific methods (e.g., XLM Conneau and Lample (2019), XLM-R Conneau et al. (2019), XLM-V Liang et al. (2023), XGLM Lin et al. (2022), PaLM-2 Anil et al. (2023), PolyLM Wei et al. (2023a), mLongT5 Uthus et al. (2023)). For instance, the series of XLM Conneau and Lample (2019); Conneau et al. (2019); Liang et al. (2023) proposes a translation language modeling (TLM) pre-training task, utilizing parallel multilingual data to improve cross-lingual model pre-training, and it extends MLM by using batches of parallel sentences instead of consecutive sentences. Moreover, the sampling strategies are important to balance the proportion of multiple languages in training data, which facilitates models to learn multilingual representation Lin et al. (2022); Anil et al. (2023); Chung et al. (2023). PoLyLM Wei et al. (2023a) integrates bilingual alignment within the training corpus and implements a curriculum learning strategy to maintain a balanced representation of various languages during the pre-training phase. The model proposes a multilingual self-instruct approach, enabling the automatic generation of diverse multilingual instructions for fine-tuning the model.
An inevitable part of the multilingual learning paradigm for PLMs is to deal with vocabulary Dabre et al. (2020); Garcia et al. (2021); Huang et al. (2022). The “out-of-vocabulary (OOV)” tokens hurts the translation performance naturally for multilingual learning Zhang et al. (2022c). To alleviate this issue, ByT5 Xue et al. (2022) is a token-free model that operates directly on raw text (bytes or characters) which simplifies the process of different languages.

### v2

@@ 2.3 Multilingual Paradigm Transition
As shown in Figure 3, multilingual pre-trained language models (PLMs) are dominated by small language models (i.e., parameters less than 7B). They are designed for specific tasks, such as multilingual neural machine translation (NMT) [264, 279, 280], multilingual question answering (QA) [281, 282, 283] and multilingual reasoning [252, 284, 285], etc. To indicate the language of target tasks, a prepending language token is appended to each source sentence [264, 278]. Take the multilingual NMT for example, formally, given the source sentence $\mathbf{x^{\prime}}=(x_{1},x_{2},...,x_{n})$, where each $x_{i}$ is a token in the source language, the modified source sentence is represented as $\mathbf{x}=(l_{i},x_{1},x_{2},...,x_{n})$. Here, $l_{i}$ denotes the target language. Correspondingly, the target sentence is represented as $\mathbf{y}=(y_{1},y_{2},...,y_{m})$, where each $y_{j}$ represents a token in the target language. The probability of a target sentence is given by:
@@
However, the current model paradigm has gradually shifted to Artificial General Intelligence (AGI) which possesses the ability like human intelligence to perform a wide range of tasks [1, 2, 286], including understanding, learning, cognition, communication and others. Unlike small language models, which are designed for specific tasks, AGI would have the capacity for generalization and adaptation to novel situations, potentially surpassing human capabilities in various domains. With the increasing scale of language models, LLMs improve the performance of downstream tasks compared to small language models [276], which is closer to the goal of AGI. For the sake of generality, the task form becomes general tasks instead of specific tasks, and prepending the language is not allowed. Hence, most existing LLMs typically do not specify whether they function as multilingual models or detail the number of languages they support. We consider that if a certain amount of multilingual data is used in the training corpus, it can be regarded as a multilingual LLM. Among the existing models, only a few are explicitly called “multilingual LLM” (e.g., BLOOM [18], BayLing [20], etc), yet there are also other models that possess multilingual capabilities [46, 21]. To distinguish LLMs from multilingual LLMs, we propose a definition of “LLMs with multilingualism”.
@@ 3 Large Language Models with Multilingual Capability
GPT-3 [30] 20-05 13B, 175B OpenAI GPT-2 Closed -
mT5 [31] 20-10 13B Google T5 Open 101
ByT5 [32] 21-05 13B Google T5 Open 101
Gopher [33] 21-12 280B DeepMind - Open -
LaMDA [34] 22-01 137B Google - Open -
OPT [35] 22-04 175B Meta - Open -
PaLM [36] 22-04 8B, 62B, 540B Google - Closed 124
mGPT [37] 22-04 13B - GPT-3 Open 61
BLOOM [18] 22-07 176B BigScience - Open 46
FLAN-T5 [56] 22-10 11B Google T5 Open 60
FLAN-PaLM [56] 22-10 8B, 62B, 540B Google PaLM Open 60
ChatGLM [57] 22-10 130B ZHIPU GLM Open 2
GPT-3.5 [1] 22-11 - OpenAI GPT Closed -
XGLM [38] 22-11 7.5B Meta AI - Open 30
LLaMA [3] 23-02 7B, 13B, 33B, 65B Meta - Open -
GPT-4 [2] 23-03 - OpenAI - Closed -
Alpaca [58] 23-03 7B StandFord LLaMA Open -
PANGU-$\sum$ [39] 23-03 1085B Huawei - Closed 26
Pythia [40] 23-04 12B EleutherAI - Open -
ParroT [59] 23-04 7B Tencent Alpaca Open -
BigTrans [17] 23-05 13B - LLaMA Open 102
PaLM-2 [41] 23-05 340B Google PaLM Closed -
Vicuna [60] 23-06 13B LMSYS LLaMA Open -
InternLM [42] 23-06 104B Shanghai AI Laboratory - Open -
BayLing [20] 23-06 7B, 13B ICT/CAS - Open -
PolyLM [43] 23-07 13B DAMO Academy - Open 18
LLaMA-2 [3] 23-07 7B, 13B, 34B, 70B Meta LLaMA Open -
Baichuan-2 [44] 23-09 7B, 13B Baichuan - Open -
Qwen [45] 23-09 7B, 14B Alibaba - Open -
OpenChat [21] 23-09 7B, 13B Tsinghua LLaMA Open -
Mistral [46] 23-10 7B Mistral - Open -
Gemini [47] 23-12 - DeepMind - Closed 40+
TigerBot [48] 23-12 7B, 13B, 70B, 180B Tiger - Open -
YAYI-2 [49] 23-12 30B IACAS - Open -
DeepSeek [50] 24-01 67B DeepSeek AI - Open -
GLM-4 [61] 24-01 - ZHIPU GLM Close -
Orion [51] 24-01 14B OrionStar - Open 8+
TeleChat [52] 24-01 7B, 12B China Telecom - Open -
Aya [62] 24-02 13B Cohere mT5 Open 101
Claude3 [53] 24-03 - Anthropic - Closed -
InternLM-2 [54] 24-03 7B, 20B Shanghai AI Lab Open -
LLaMA-3 [55] 24-04 8B, 70B Meta - Open 30+
Phi-3 [52] 24-04 7B, 14B Microsoft - Open -
@@ 3.1 Training from Scratch
To obtain a language model with multilingual capability, a common practice is to leverage all available data in different languages for training. The language sampling algorithms are usually applied to control the importance of each language. As shown in Figure 3, multilingual pre-trained language PLMs before 2022 mainly consisted of two structures: (1) encoder-only designed for natural language understanding tasks and (2) encoder-decoder designed for natural language generation tasks. The parameters of these PLMs are not comparable with the existing LLMs. Recent studies show that according to the scaling law [276, 287, 288], the scale of the parameter has significant impacts on the performance of the models, i.e., the larger models lead to better performance. The researchers also observe that when the parameter scale of the language models exceeds a certain level, they not only achieve significant performance improvements but also find some special capabilities not observed within small-scale PLMs [276, 287]. To distinguish the difference in parameter scale, the research community specifies the term “LLMs” compared with PLMs based on a significant scale [289].
@@
In addition to constructing multilingual data in the pre-training stage, the cross-lingual capabilities can also be improved by some specific methods (e.g., XLM [278], XLM-R [292], XLM-V [293], XGLM [38], PaLM-2 [41], PolyLM [43], mLongT5 [291]). For instance, the series of XLM [278, 292, 293] proposes a translation language modeling (TLM) pre-training task, utilizing parallel multilingual data to improve cross-lingual model pre-training, and it extends MLM by using batches of parallel sentences instead of consecutive sentences. Moreover, the sampling strategies are important to balance the proportion of multiple languages in training data, which facilitates models to learn multilingual representation [38, 41, 294]. PoLyLM [43] integrates bilingual alignment within the training corpus and implements a curriculum learning strategy to maintain a balanced representation of various languages during the pre-training phase. The model proposes a multilingual self-instruct approach, enabling the automatic generation of diverse multilingual instructions for fine-tuning the model.
An inevitable part of the multilingual learning paradigm for PLMs is to deal with vocabulary [295, 296, 297]. The “out-of-vocabulary (OOV)” tokens hurts the translation performance naturally for multilingual learning [298]. To alleviate this issue, ByT5 [32] is a token-free model that operates directly on raw text (bytes or characters) which simplifies the process of different languages.